A skateboarder rolls horizontally off the top of a staircase and lands at the bottom. The staircase has a horizontal length of 12.0\,\text m12.0m12, point, 0, start text, m, end text with a vertical drop of 2.00\,\text m2.00m2, point, 00, start text, m, end text. We can ignore air resistance.
1
Expert's answer
2020-11-27T00:40:51-0500
"t=\sqrt{\frac{2h}{g}}\\v=\frac{x}{t}=x\sqrt{\frac{g}{2h}}=12\sqrt{\frac{9.8}{2(2)}}=18.8\ m"